[Hallc_running] [New Logentry] Last Pion-LT Hall C RC summary: Friday, September 9, 2022

gaskelld at jlab.org gaskelld at jlab.org
Fri Sep 9 08:45:02 EDT 2022

Logentry Text:
Another productive day yesterday. In the afternoon, Dave Mack worked with MCC to perform a low current BCM calibration using the Faraday Cup (https://logbooks.jlab.org/entry/4039232) - thanks to Hall B and D for giving us the time for this invasive procedure.  After the BCM calibration, we took data for the Q2=3.85, W=2.62, low epsilon SHMS center (8 degrees) setting for pi- on dummy and LD2.  We stayed at this setting until 8 am this morning, accumulating a total of 5156 good pion events on LD2.

Just before 8 am, we requested beam off and Dave Meekins started the target warmup.  SHMS and HMS magnets are ramped to zero and the power supplies turned off.  Radiation survey of the hall is underway, and we will go to Restricted Access when complete.

There will be no RC meeting today.

This marks the end of the pion-LT running in Hall C.  Thanks to everyone for their help and support over the course of the long run of this experiment!


This is a plain text email for clients that cannot display HTML.  The full logentry can be found online at https://logbooks.jlab.org/entry/4039524
-------------- next part --------------
An HTML attachment was scrubbed...
URL: <https://mailman.jlab.org/pipermail/hallc_running/attachments/20220909/58a6f90e/attachment.html>

More information about the Hallc_running mailing list